use Cramer's Rule to find the solution of the systems of linear equation in terms of the parameter K.


2x - 3y = K

x + 2y = -2


1
Expert's answer
2021-04-14T13:11:43-0400

"\Delta = \left| {\begin{matrix}
2&{ - 3}\\
1&2
\end{matrix}} \right| = 4 + 3 = 7"

"{\Delta _1} = \left| {\begin{matrix}
K&{ - 3}\\
{ - 2}&2
\end{matrix}} \right| = 2K - 6"

"{\Delta _2} = \left| {\begin{matrix}
2&K\\
1&{ - 2}
\end{matrix}} \right| = - 4 - K"

Then

"x = \frac{{{\Delta _1}}}{\Delta } = \frac{{2K - 6}}{7}"

"y = \frac{{{\Delta _2}}}{\Delta } = \frac{{ - 4 - K}}{7}"

Answer: "x = \frac{{2K - 6}}{7}, y = \frac{{ - 4 - K}}{7}"